OpenStudy (anonymous):

A tower in Bloomington sends radio signal a certain distance (in miles) according to the equation: x2 + y2 = 6,400. The Richards family lives 50 miles north and 70 miles west of Bloomington. Can the Richards family receive the radio signal from Bloomington? In complete sentences, explain your reasoning.

OpenStudy (anonymous):

last question i need to complete assignment please help :)

OpenStudy (anonymous):

please help

ganeshie8 (ganeshie8):

do you have circles & their equations in ur syllabus ?

OpenStudy (anonymous):

i have the circle and equations for them yes i just dont no how to do this unless its like x^2+Y^2+50x+70y=6400

ganeshie8 (ganeshie8):

ok. below equation also represents a circle with radius r x^2 + y^2 = r^2

OpenStudy (anonymous):

yes i have that but where do the numbers 50 and 70 come in

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

say that dot is the tower in bloomingtom

OpenStudy (anonymous):

okay

ganeshie8 (ganeshie8):

it is sending radio signals whose range is represented by the equation x^2+y^2 = 80^2

ganeshie8 (ganeshie8):